# CRABS: A syntactic-semantic pincer strategy for bounding LLM interpretation of Python notebooks

### v1

@@ 4.2 Semantic-aware Phase
We hypothesize that pre-trained language models can help resolve ambiguous inter-cell I/O sets identified by the syntactic phase. We resolve two different kinds of ambiguities, (1) checking if data is an input of a given code cell and (2) checking if data is an output candidate of a given code cell. An LLM resolves these ambiguities cell by cell using a zero-shot in-context learning approach. Example prompts are included in Figure 16 and Figure 17 of the Appendix. Requesting binary answers (i.e., “yes” or “no”) eliminates hallucinations of variables; all data inputs and output candidates are selected from the given syntactic phase results. Prompts include information about shared references defined in previous cells only when ambiguous data is among the shared references. Since these statements and variable names for shared references are extracted cell by cell in sequence in the syntactic phase, LLMs are prompted for each code cell independently, which helps improve time efficiency.
@@ 5.1 Dataset
We determined which notebooks met these criteria by carefully inspecting their contents. We inspected the 104 most up-voted Python notebooks to find 50 meeting these criteria. Reasons for excluding the other 54 notebooks are indicated in Table 4 of the Appendix. For the ones we selected, we extracted corresponding kernel version IDs from the Meta Kaggle dataset (Risdal & Bozsolik, 2022), and retrieved these notebooks from the Meta Kaggle Code dataset (Plotts & Risdal, 2023).
@@ 5.3 Main Results
All experiments were conducted using the GPT-4o (the “gpt-4o-2024-08-06” version) model with the temperature set to 0.
We compare the performance of CRABS against a baseline approach which employs zero-shot prompting of an LLM for the whole notebook with carefully designed chain-of-thought prompts (Wei et al., 2022). An example prompt is shown (in part) in Figure 15 of the Appendix. We observe that the LLM fails to understand 10 out of 50 (20%) notebooks, returning the wrong number of cells and earning zero values for all metrics. These 10 notebooks are long with 52 cells on average, where the LLM faces long-context challenges. For other notebooks, the LLM sometimes hallucinates, describing variables not present at all in the reported cells.
Table 2 summarizes the performance of CRABS for the notebook understanding task. These results are average scores across 50 notebooks in our dataset. For cell execution dependency graphs, the precision for lower estimate and the recall for upper estimate are 100%, which are consistent with the “subgraph/supergraph” relations. Metrics for lower estimate and upper estimate are not intended for evaluation, but for reflecting the characteristics of our notebooks. Compared with the baseline, CRABS yields non-zero scores for all notebooks using the cell-by-cell prompting strategy. Table 2 shows a percentage-point increase of 28 in F1 score, 32 in accuracy, and 48 in EM for information flow graphs; and percentage-point increases of 26 in F1 score, 30 in accuracy, and 46 in EM for cell execution dependency graphs. These results demonstrate the effectiveness of CRABS for the notebook understanding task.
37 information flow graphs extracted using CRABS exactly matches the ground truth. Within these 37 exactly matched notebooks, 4 notebooks return an identical lower and upper estimate. As an LLM resolves ambiguities between these estimates, the LLM is not used for these 4 notebooks. Therefore, metrics in Table 2 cannot accurately reflect the true capability of an LLM for resolving ambiguities. We evaluate such capability at a finer level of granularity, and find that 1397 out of 1425 (98%) ambiguities are correctly resolved by an LLM.
@@
5.4 Ablation Study
Because cell execution dependency graphs are extracted from information flow graphs, we only discuss the information flow graphs in this section. For completeness, the results for cell execution dependency graphs are reported in Table 7 and Table 8 of the Appendix. Detailed results for each notebook are reported in Table 9 of the Appendix.
A1. Effectiveness of the syntactic phase. Instead of using the syntactic-semantic hybrid approach, we test the performance of an approach without the syntactic phase, that asks an LLM to identify cell inputs and output candidates without providing any hints from the syntactic processor. An example prompt is in Figure 18 of the Appendix. Table 3 shows percentage-point decreases of 7 in F1 score, 10 in accuracy, and 36 in EM, demonstrating the effectiveness of syntactic parsing.
A2. Effectiveness of cell-by-cell prompting. Instead of using the cell-by-cell prompting approach, we test the performance of using a single prompt for the whole notebook. Part of the example prompt is shown in Figure 19 of the Appendix. Table 3 shows percentage-point decreases of 16 in F1 score, 20 in accuracy, and 38 in EM. The main reason it drops dramatically is the long-context challenge, which makes LLMs fail to handle multiple code cells using a single prompt. Five notebooks, each containing an average of 54 code cells, return the wrong number of cells. Metrics over the remaining 45 notebooks are shown in Table 6 of the Appendix, which demonstrates that the single prompt approach performs worse than cell-by-cell prompting.
@@ Appendix A Appendix
A.1 An example case depending on runtime information
Figure 10 provides an example when some inter-cell data flows only exisit in certain executions. Cell 1 loads the phone_data, cell 2 samples the phone_data if the number of rows is greater than 100, and cell 3 previews the phone_data. Line 3 in cell 2 will be executed only when the phone_records.csv file has more than 100 rows. YesWorkflow captures possible data flows between program blocks, including cases when the line is executed or skipped.
Figure 10: An example case depending on runtime information.
A.2 Syntactic Parsing
A.2.1 Distinct data and code names
@@
Figure 11 shows an example case that violates the assumption of distinct data and code names. Cell 1 defines a variable named add_one, cell 2 reuses the variable name as a function name, and cell 3 calls the function. Our system cannot identify that the add_one data is a data output candidate of cell 1 and a data input of cell 2, which further fails to extract the data flow (cell 1, cell 2, add_one) and the direct dependency (cell 2, cell 1). However, the code flow (cell 2, cell 3, add_one) and the direct dependency (cell 3, cell 2) are still captured.
Figure 11: An example case for reusing a data name as a code name.
A.2.2 Simplified function parsing approach
@@
Figure 12 shows an example case of the limitation of the simplified function parsing approach. Cell 1 defines a global variable data_file_path and cell 2 uses it by calling the load_data function without passing any argument. Therefore, our system cannot identify that data_file_path is as a data input of cell 2, which further fails to recognize the information flow (cell 1, cell 2, data_file_path) and the direct dependency (cell 2, cell 1).
Figure 12: An example case for the simplified function parsing approach.
A.2.3 Mixed code issue
@@
Figure 13 shows an example case for the mixed code issue. Cell 1 uses the %%capture magic command to capture the standard output of the code cell, storing it as a variable named captured_stdout that is later used by cell 2. However, our system skips cell 1 due to the magic command and cannot identify captured_stdout as a data output candidate of cell 1, which fails to extract the information flow (cell 1, cell 2, captured_stdout) and the direct dependency (cell 2, cell 1).
Figure 13: An example case for the mixed code issue.
A.2.4 Conditional statements in a loop
@@
Figure 14 shows an example case for conditional statements in a loop. Cell 1 defines count and cell 2 have a if-statement inside a loop, which assigns 0 to count at the first iteration and increments count by one during the later two iterations. There are two possible execution orders, i.e., line 4 first, or line 6 first. If line 4 is executed earlier than line 6, then the count used in cell 2 is from itself, otherwise it is from cell 1. The lower estimate considers the intersection of all possible branch execution orders, finding cell 1 and cell 2 are independent. However, the upper estimate considers the union of all possible branch execution orders, resulting in an information flow (cell1, cell2, count) and a direct dependency (cell 2, cell 1).
Figure 14: An example case for conditional statements in a loop.
@@
06 ldfreeman3/a-data-science-framework-to-achieve-99-accuracy 35 $1$ 103 146
@@
Figure 18: An example prompt for the ablation study A1 (i.e., w/o syntactic phase).
Figure 19: Part of an example prompt for the ablation study A2 (i.e., w/o cell-by-cell prompting). The real prompt includes all code cells.
@@
baseline CRABS A1 A2 baseline CRABS A1 A2
@@
Table 9: Detailed results for notebook understanding task. The A1 and A2 columns denote ablation study A1 (i.e., w/o syntactic phase) and ablation study A2 (i.e., w/o cell-by-cell prompting). Values indicate the F1 score of a notebook, where the notebook can be referred to via the ID column and Table 5.

### v2

@@ 4.2 Semantic-aware Phase
We hypothesize that pre-trained language models can help resolve ambiguous members of the inter-cell I/O set identified by the syntactic phase. We resolve two different kinds of ambiguities, (1) checking if data is an input of a given code cell and (2) checking if data is an output candidate of a given code cell. An LLM resolves these ambiguities cell by cell using a zero-shot in-context learning approach. Example prompts are included in Figure 15 and Figure 16 of the Appendix. Requesting binary answers (i.e., “yes” or “no”) eliminates hallucinations of variables; all data inputs and output candidates are selected from the given syntactic phase results. Prompts include information about shared references defined in previous cells only when shared references are among the ambiguous data. Since the names of shared references are identified in the syntactic phase, the LLM can be prompted for each code cell independently, improving time efficiency.
@@ 5.1 Dataset
We determined which notebooks met these criteria by carefully inspecting their contents. We inspected the 104 most up-voted Python notebooks to find 50 meeting these criteria. Reasons for excluding the other 54 notebooks are indicated in Table 5 of the Appendix. For each notebook we selected, we extracted the corresponding kernel version ID from the Meta Kaggle dataset (Risdal & Bozsolik, 2022), and retrieved the notebook from the Meta Kaggle Code dataset (Plotts & Risdal, 2023).
@@ 5.3 Main Results
Results were obtained using the gpt-4o-2024-08-06 version of the GPT-4o model (OpenAI, 2024a) with temperature set to 0. We compared the performance of CRABS against a baseline approach which employs zero-shot prompting of an LLM for the whole notebook with carefully designed chain-of-thought prompts (Wei et al., 2022). An example prompt is shown (in part) in Figure 14 of the Appendix. We observe that the LLM completely fails to understand 10 out of 50 (20%) notebooks, returning the wrong number of cells and earning zero values for all metrics. These 10 notebooks are long with 52 cells on average, where the LLM faces long-context challenges. For other notebooks, the LLM sometimes hallucinates, describing variables not present at all in the reported cells.
Table 2 summarizes the performance of CRABS for the notebook understanding task. These results are average scores across 50 notebooks in our dataset. We observe a percentage-point increase of 28 in F1 score, 32 in accuracy, and 48 in EM for information flow graphs; and percentage-point increases of 26 in F1 score, 30 in accuracy, and 46 in EM for cell execution dependency graphs. These results demonstrate the effectiveness of CRABS for the notebook understanding task. We include in this table metrics for the lower estimate and upper estimate components of CRABS for comparison with the full CRABS approach. For the transitive dependencies, the precision for lower estimate and the recall for upper estimate are each 100%, consistent with the subgraph and supergraph relations described in Section 4.1. We note that in contrast to the baseline, CRABS yields plausible outputs (i.e., earns non-zero contributions for each metric) for all notebooks; we attribute this improvement to the cell-by-cell prompting strategy CRABS employs which avoids the long-context challenges faced by the baseline method.
CRABS yields information flow graphs that exactly match the ground truth for 37 notebooks. Of these, 4 return an identical lower and upper estimate. As an LLM resolves ambiguities between these estimates, the LLM is not used for these 4 notebooks. The metrics in Table 2 thus do not fully reflect the capability of the LLM for resolving ambiguities. We evaluate such capability at a finer level of granularity, finding that 1397 out of 1425 (98%) ambiguities are correctly resolved by the LLM.
@@
5.4 Ablation Studies
Because cell execution dependency graphs are extracted from information flow graphs, we only discuss the information flow graphs in this section. For completeness, the results of the ablation studies with respect to cell execution dependency graphs are reported in Table 8 and Table 9 of the Appendix. Detailed results for each notebook are reported in Table 10 of the Appendix.
S1. Effectiveness of the syntactic phase. Instead of using the syntactic-semantic hybrid approach, we test the performance of an approach without the syntactic phase, asking the LLM to identify cell inputs and output candidates without providing any hints from the syntactic processor. An example prompt is in Figure 17 of the Appendix. Table 3 shows percentage-point decreases of 7 in F1 score, 10 in accuracy, and 36 in EM, demonstrating the effectiveness of syntactic parsing.
S2. Effectiveness of cell-by-cell prompting. Instead of using the cell-by-cell prompting approach, we test the performance of using a single prompt for the whole notebook. Part of the example prompt is shown in Figure 18 of the Appendix. Table 3 shows percentage-point decreases of 16 in F1 score, 20 in accuracy, and 38 in EM. The main reason it drops dramatically is the long-context challenge, which causes the LLM to fail to handle multiple code cells using a single prompt. Five notebooks, each containing an average of 54 code cells, return the wrong number of cells. Detailed metrics for the remaining 45 notebooks are shown in Table 7 of the Appendix. These results demonstrate that cell-by-cell prompting outperforms the single prompt approach.
@@ 6 Conclusion
7 Future Work
Work is underway to enhance CRABS to extract and represent multiple distinct information flow graphs for a single notebook. This refinement will enable more precise and comprehensive analysis by capturing multiple alternative information flows that may emerge under different conditions.
We are currently investigating approaches for optimizing the overall execution performance of CRABS, including reducing the total latency of the notebook understanding task. Preliminary analysis shows that a sequential implementation of CRABS achieves lower latency for this task than the baseline for notebooks with a small number of ambiguities, but that this latency increases beyond that of the baseline approach as the number of ambiguities grows. Further experiments demonstrate that a concurrent implementation of CRABS consistently outperforms the baseline in terms of latency, regardless of the number of ambiguities. These findings (see Appendix A.2) point to the potential for CRABS to achieve the performance required for practical, interactive applications. Smaller LLMs may offer further improvements in execution speed; see Appendix A.3 for early results evaluating CRABS on such models.
Finally, in addition to serving as an effective neuro-symbolic technique for notebook understanding, we view CRABS as the foundation for an experimental platform for investigating the information used, the skills employed, and the sub-problems that must be solved for any intelligence, artificial or natural, to understand a Python notebook via its information flow graph and the corresponding cell execution dependency graph.
@@ Appendix A Appendix
A.1 Illustrative Python notebooks
A.1.1 Notebook with information flow dependent on input data
The notebook illustrated in Figure 7 exhibits different patterns of information flow depending on the input dataset. Because line 3 in cell 2 will be executed only when the phone_records.csv file has more than 100 rows, information flows from cell 2 to cell 3 only for runs on such input files.
Figure 7:
A.1.2 Notebook with clashing variable and function names
@@
Figure 8 shows an example case that violates the assumption of distinct data and code names. Cell 1 defines a variable named add_one, cell 2 reuses the variable name as a function name, and cell 3 calls the function. Our system cannot identify that the add_one variable is a data output candidate of cell 1 and a data input of cell 2, resulting in a further failure to extract the data flow (cell 1, cell 2, add_one) and the cell execution dependency (cell 2, cell 1).
Figure 8:
A.1.3 Global variable accessed in a function
@@
Cell 1 of the notebook in Figure 9 defines a global variable data_file_path and cell 2 accesses it directly from the load_data function. Our system cannot identify that data_file_path is a data input of cell 2, resulting in a further failure to recognize the information flow (cell 1, cell 2, data_file_path) and the cell execution dependency (cell 2, cell 1).
Figure 9:
A.1.4 Notebook with non-Python statements
@@
Cell 1 in Figure 10 uses the %%capture magic command to capture the standard output of the code cell, storing it as a variable named captured_stdout that is later used by cell 2. Our system skips cell 1 due to the magic command, and fails to extract the information flow (cell 1, cell 2, captured_stdout) and the cell execution dependency (cell 2, cell 1).
Figure 10:
A.1.5 Conditional statements in a loop
@@
Cell 1 in Figure 11 defines count and cell 2 has an if statement inside a loop, assigning 0 to count on the first iteration and incrementing count by one during each of the later two iterations. While the lower estimate considers the intersection of all possible branch execution orders, finding cell 1 and cell 2 to be independent; the upper estimate considers the union of all possible branch execution orders, finding the information flow (cell1, cell2, count) and a cell execution dependency (cell 2, cell 1).
Figure 11:
A.2 Execution time analysis
Figure 12 shows that, when using the sequential implementation of CRABS, the execution time for a notebook, defined as the total of the latencies of the individual requests involved in analyzing the notebook, is proportional to the total number of ambiguities. The ordinary least squares (OLS) linear regression yields a slope of 0.94 seconds per ambiguity and an $R^{2}$ of 0.96. In contrast, the OLS fit of the baseline execution time against the number of ambiguities is substantially weaker with an $R^{2}$ of 0.54, though its slope of 0.22 seconds per ambiguity remains informative. The intersection point of these two regression lines supports the observation that CRABS is generally up to 5 times faster than the baseline for notebooks with fewer than 14 ambiguities, and as much as 7 times slower than the baseline when there are more ambiguities. A separate regression against the total number of information flows in a notebook (Figure 13) better predicts the baseline execution time with an $R^{2}$ of 0.76. For the 4 notebooks containing no ambiguities (i.e., for each cell in these notebooks the parser-generated lower and upper estimate of the inter-cell I/O sets are identical), CRABS is approximately 3 orders of magnitude faster than the baseline because no prompting of an LLM is necessary; these 4 notebooks were excluded from the linear regressions discussed above.
While the sequential implementation of CRABS resolves exactly one ambiguity per request in a sequential manner, the concurrent implementation sends requests to an LLM without waiting for prior responses and achieves performance improvements ranging from 4 to 24 times faster than the baseline on a per-notebook basis, with per-notebook execution times ranging from 0.6 to 8.3 seconds.
Figure 12: Scatter plot of execution time vs. number of ambiguities per notebook with least squares lines.
Figure 13: Scatter plot of execution time vs. number of information flows per notebook with least squares lines.
A.3 Effectiveness of CRABS on LLMs other than GPT-4o
Table 4 validates the effectiveness of CRABS on smaller models, including open- and closed-source, general-purpose and code-focused language models—GPT-4o mini (OpenAI, 2024b), Qwen3-8B (Yang et al., 2025), and Qwen2.5-Coder-7B-Instruct (Hui et al., 2024). We observe that smaller models face greater long-context challenges. For example, the GPT-4o mini model fails to understand 25 notebooks under the baseline setting and 20 notebooks without cell-by-cell prompting, returning the wrong number of cells and resulting in zero values for all metrics in each of these notebooks.
LLM Method Information Flows (%)
Prec. Rec. F1 Acc. EM
GPT-4o-mini baseline 43.19 40.65 41.53 37.40 12
(2024-07-18) CRABS 83.01 82.87 82.94 74.85 38
w/o syntactic phase 71.58 65.16 67.73 55.09 10
w/o cell-by-cell prompting 54.69 53.57 54.01 50.55 28
Qwen3-8B baseline 45.75 38.92 41.36 34.77 6
CRABS 75.04 73.75 74.31 63.48 22
w/o syntactic phase 87.56 45.77 57.21 43.78 8
w/o cell-by-cell prompting 37.07 36.86 36.95 32.29 16
Qwen2.5-Coder baseline 19.47 13.87 15.80 12.64 2
(7B-Instruct) CRABS 77.13 76.01 76.52 66.21 26
w/o syntactic phase 64.90 56.64 59.81 47.24 6
w/o cell-by-cell prompting 35.55 33.49 34.36 31.71 16
Table 4: CRABS results for LLMs other than GPT-4o.
A.4 Additional figures and tables
@@
06 ldfreeman3/a-data-science-framework-to-achieve-99-accuracy 35 1111 103 146
@@
Figure 17: An example prompt for the ablation study S1 (i.e., w/o syntactic phase).
Figure 18: Part of an example prompt for the ablation study S2 (i.e., w/o cell-by-cell prompting). The real prompt includes all code cells.
@@
baseline CRABS S1 S2 baseline CRABS S1 S2
@@
Table 10: Detailed results for notebook understanding task. The S1 and S2 columns denote ablation study S1 (i.e., w/o syntactic phase) and ablation study S2 (i.e., w/o cell-by-cell prompting). Values indicate the F1 score of a notebook, where the notebook can be referred to via the ID column and Table 6.